arectangular coil of areaA,no of turnsN isrotated of f rev per sec in a uniform magneticfieldB,the field being perpendicular to the coil.prove that themaximum emf induced in the coil is 2(pi)fNBA.


 

Grade:

1 Answers

gauhar singh
11 years ago

Let at any instant angle between flux n perpendicular to face of coil be   Wt where W is angular velocity and time is t.

flux(F) through coil is ABcosWt i.e. dot product of field strenght n aerial vector of face of coil

d(F)/dt= induced emf by faraday''s law

on difff,

induced emf (E) = -ABWSinWt

But W=2(pi)f

hence E = -2(pi)fABSinWt . max value of sinWt =1 hence proved